A sample of Helium occupies 200 L at 30ºC under a pressure of 842 torr. What is the volume it occupies at standard temprature and pressure (STP)?
Answer #1

Standard temperature = 273 K = T2

Standard pressure = 1 atm = 760 torr = P2

Initial temperature T1 = 273 + 30 = 303 K

Initial pressure P1 = 842 torr

Initial volume V1 = 200 L

Let volume at STP = V2

Using combine gas law, P2V2/T2 = P1V1/T1

V2 = (P1V1/T1) ×T2/P2 = 199.64 L

Volume occupied at STP = 199.64 L